Determine all eigenvalues and the corresponding eigenspaces for the matrix 𝐴 = [ −9 4 4 −8 3 4 −16 8 7 ]


1
Expert's answer
2021-09-12T23:52:10-0400
"A-\lambda I=\begin{pmatrix}
 -9-\lambda & 4 & 4 \\
 -8& 3-\lambda & 4 \\
 -16 & 8 & 7-\lambda \\
\end{pmatrix}"

"\det(A-\lambda I)=\begin{vmatrix}
 -9-\lambda & 4 & 4 \\
 -8& 3-\lambda & 4 \\
 -16 & 8 & 7-\lambda \\
\end{vmatrix}"


"=(-9-\lambda)\begin{vmatrix}
 3-\lambda & 4 \\
 8 & 7-\lambda
\end{vmatrix}-4\begin{vmatrix}
 -8 & 4 \\
 -16 & 7-\lambda
\end{vmatrix}"

"+4\begin{vmatrix}
 -8 & 3-\lambda \\
 -16 & 8
\end{vmatrix}=(-9-\lambda)(21-10\lambda+\lambda^2-32)"

"-4(-56+8\lambda+64)+4(-64+48-16\lambda)"

"=99+11\lambda+90\lambda+10\lambda^2-9\lambda^2-\lambda^3"

"-32-32\lambda-64-64\lambda"

"=-\lambda^3+\lambda^2+5\lambda+3=0"

"-\lambda^2(\lambda+1)+2\lambda(\lambda+1)+3(\lambda+1)=0"

"-(\lambda+1)(\lambda^2-2\lambda-3)=0"

"-(\lambda+1)^2(\lambda-3)=0"



"\lambda_1=-1, \lambda_2=-1, \lambda_3=3"


These are the eigenvalues: "-1, -1, 3."


"\lambda=-1"


"A-\lambda I=\begin{pmatrix}
 -9+1 & 4 & 4 \\
 -8& 3+1 & 4 \\
 -16 & 8 & 7+1 \\
\end{pmatrix}"

"=\begin{pmatrix}
 -8 & 4 & 4 \\
 -8 & 4 & 4 \\
 -16 & 8 & 8 \\
\end{pmatrix}"

"R_2=R_2-R_1"


"\begin{pmatrix}
 -8 & 4 & 4 \\
 0 & 0 & 0 \\
 -16 & 8 & 8 \\
\end{pmatrix}"

"R_3=R_3-2R_1"


"\begin{pmatrix}
 -8 & 4 & 4 \\
 0 & 0 & 0 \\
 0 & 0 &0 \\
\end{pmatrix}"

"R_1=R_1\/(-8)"


"\begin{pmatrix}
 1 & -1/2 & -1/2 \\
 0 & 0 & 0 \\
 0 & 0 &0 \\
\end{pmatrix}"

"\begin{pmatrix}
 1 & -1/2 & -1/2 \\
 0 & 0 & 0 \\
 0 & 0 &0 \\
\end{pmatrix}\begin{pmatrix}
 x_1 \\
x_2 \\
x_3 \\
\end{pmatrix}=\begin{pmatrix}
0 \\
0 \\
0 \\
\end{pmatrix}"

If we take "x_2=t, x_3=s," then "x_1=\dfrac{1}{2}t+\dfrac{1}{2}s"

Thus


"\vec x=\begin{pmatrix}
s/2+t/2 \\
1 \\
1 \\
\end{pmatrix}s=\begin{pmatrix}
1/2 \\
1 \\
0 \\
\end{pmatrix}t+\begin{pmatrix}
1/2 \\
0 \\
1 \\
\end{pmatrix}s"

The eigenvectors are


"\begin{pmatrix}
1/2 \\
1 \\
0 \\
\end{pmatrix},\begin{pmatrix}
1/2 \\
0 \\
1 \\
\end{pmatrix}"

"\lambda=3"

"A-\lambda I=\begin{pmatrix}
 -9-3 & 4 & 4 \\
 -8& 3-3 & 4 \\
 -16 & 8 & 7-3 \\
\end{pmatrix}"

"=\begin{pmatrix}
 -12 & 4 & 4 \\
 -8 & 0 & 4 \\
 -16 & 8 & 4 \\
\end{pmatrix}"

"R_2=R_2-2R_1\/3"

"\begin{pmatrix}
 -12 & 4 & 4 \\
 0 & -8/3 & 4/3 \\
 -16 & 8 & 4 \\
\end{pmatrix}"

"R_3=R_3-4R_1\/3"

"\begin{pmatrix}
 -12 & 4 & 4 \\
 0 & -8/3 & 4/3 \\
 0 & 8/3 & -4/3 \\
\end{pmatrix}"

"R_3=R_3+R_2"

"\begin{pmatrix}
 -12 & 4 & 4 \\
 0 & -8/3 & 4/3 \\
 0 & 0 & 0 \\
\end{pmatrix}"

"R_2=-3R_2\/8"


"\begin{pmatrix}
 -12 & 4 & 4 \\
 0 & 1 & -1/2 \\
 0 & 0 & 0 \\
\end{pmatrix}"

"R_1=-R_1\/12"


"\begin{pmatrix}
 1 &-1/3 & -1/3 \\
 0 & 1 & -1/2 \\
 0 & 0 & 0 \\
\end{pmatrix}"

"R_1=R_1+R_2\/3"


"\begin{pmatrix}
 1 &0 & -1/2 \\
 0 & 1 & -1/2 \\
 0 & 0 & 0 \\
\end{pmatrix}"


"\begin{pmatrix}
 1 &0 & -1/2 \\
 0 & 1 & -1/2 \\
 0 & 0 & 0 \\
\end{pmatrix}\begin{pmatrix}
 x_1 \\
x_2 \\
x_3 \\
\end{pmatrix}=\begin{pmatrix}
0 \\
0 \\
0 \\
\end{pmatrix}"

If we take "x_3=t," then "x_1=\dfrac{1}{2}t, x_2=\dfrac{1}{2}t"

Thus


"\vec x=\begin{pmatrix}
t/2 \\
t/2 \\
t \\
\end{pmatrix}=\begin{pmatrix}
1/2 \\
1/2 \\
1 \\
\end{pmatrix}t"


The eigenvector is


"\begin{pmatrix}
1/2 \\
1/2 \\
1 \\
\end{pmatrix}"



Eigenvalue: −1, multiplicity: 2, eigenvectors: "\begin{pmatrix}
1/2 \\
1 \\
0 \\
\end{pmatrix},\begin{pmatrix}
1/2 \\
0 \\
1 \\
\end{pmatrix}"


Eigenvalue: 3, multiplicity: 1, eigenvector: "\begin{pmatrix}
1/2 \\
1/2 \\
1 \\
\end{pmatrix}"